OpenStudy (anonymous):

How do you solve 2.1x + 45.2 = 3.2 - 8.4x

OpenStudy (vijay):

x(2.1 + 8.4 ) = 3.2 - 45.2 10.5 x = -42 x =-4

OpenStudy (anonymous):

how did you get the answer though

OpenStudy (anonymous):

all you have to do is add coefficents of the same variiable.

OpenStudy (anonymous):

so add 8.4x to both sides and you get 10.5x+45.2=3.2

OpenStudy (anonymous):

oops wait no i mean 10.5=-42

OpenStudy (anonymous):

10.5 as a fraction is 21/2

OpenStudy (anonymous):

oops again so sorry. 10.5x=-42 so 21/2x=-42 multiply both sides by 2 21x=-84 and divide both sides by 21. x=-4

OpenStudy (anonymous):

is that good?

OpenStudy (anonymous):

thanks!